# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ Appendix F Extended results
F.1 Training and evaluation.
Five splits are generated for each dataset, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively; the exception is ZINC, for which the predefined splits are used [27]. If available, this work considers the results reported in TopoBench for the baseline models. In the case of the HOPSE architecture and SaNN baseline, the benchmarking of the corresponding architectures follows the predefined TopoBench suite.
@@
F.2 Hyperparameter search
@@
  • Encoder dropout: {0.25, 0.5}
@@
F.3 Datasets
The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35]. For regression, the ZINC dataset is used [27]. Following standard practice in the TDL literature [42], liftings are applied to transform these graph datasets into higher-order domains. Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
Additionally, the MANTRA dataset comprises different triangulations of manifolds. A triangulation of a manifold ${\mathcal{M}}$ is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between ${\mathcal{M}}$ and the geometric realization of $K$. Concretely, triangulations encode higher-order topological information that the underlying graph, or 00-skeleton, does not encode. Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions. This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties. Thus, it aims to assess how much the model encodes topological information beyond the base graph. In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones.
Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al. [2] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number. The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT.
F.4 Runtime per epoch
In these experiments, the runtime per epoch is shown for each dataset. Table 4 shows the average runtime per epoch on the benchmarked datasets. Our method has the lowest runtime per epoch across all datasets among HOMP methods. These results validate the theoretical findings presented. In addition, HOPSE-G is faster than HOPSE-M in all benchmarks except NCI109. This aligns with complexity and scales linearly with the number $K$ of positional encodings considered.
Compared to graph message passing methods, such as GCN, GAT, and GIN, the HOPSE framework performs competitively on smaller datasets such as MUTAG, PROTEINS, NCI1, and NCI109. On larger datasets, it speeds up up to 4x in ZINC and closer to 2x in MANTRA.
Table 4: Runtime per epoch in seconds. Results are shown as mean and standard deviation. The best result per domain and per model is in bold and results within one standard deviation of the lowest per domain are highlighted in blue. The HOPSE framework is more efficient than all other HOMP-based methods across all datasets. In the larger datasets such as ZINC and MANTRA, it is more efficient than traditional message passing over graphs. The variation between the simplicial and cell domain for both models is relatively low
@@ F.4 Runtime per epoch
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 0.03 $\pm$ 0.00 0.05 $\pm$ 0.00 0.26 $\pm$ 0.01 0.27 $\pm$ 0.02 1.24 $\pm$ 0.01 6.24 $\pm$ 0.11 6.24 $\pm$ 0.10 6.43 $\pm$ 0.08 GAT 0.04 $\pm$ 0.00 0.07 $\pm$ 0.00 0.34 $\pm$ 0.02 0.33 $\pm$ 0.02 1.24 $\pm$ 0.01 6.38 $\pm$ 0.14 6.16 $\pm$ 0.12 6.44 $\pm$ 0.15 GIN 0.03 $\pm$ 0.00 0.06 $\pm$ 0.00 0.27 $\pm$ 0.02 0.27 $\pm$ 0.02 1.19 $\pm$ 0.00 6.37 $\pm$ 0.12 6.18 $\pm$ 0.06 6.46 $\pm$ 0.23 Simplicial SCN 0.09 $\pm$ 0.00 0.42 $\pm$ 0.01 1.60 $\pm$ 0.03 1.59 $\pm$ 0.02 6.93 $\pm$ 0.09 6.83 $\pm$ 0.08 7.35 $\pm$ 0.38 6.92 $\pm$ 0.08 SCCNN 0.09 $\pm$ 0.00 0.43 $\pm$ 0.01 1.65 $\pm$ 0.02 1.66 $\pm$ 0.03 10.40 $\pm$ 0.06 8.47 $\pm$ 0.63 8.25 $\pm$ 0.12 9.24 $\pm$ 0.15 SaNN 0.03 $\pm$ 0.00 0.18 $\pm$ 0.00 0.64 $\pm$ 0.02 0.65 $\pm$ 0.02 0.63 $\pm$ 0.03 7.21 $\pm$ 0.08 7.77 $\pm$ 0.15 7.34 $\pm$ 0.10 GCCN 0.04 $\pm$ 0.00 3.06 $\pm$ 0.02 0.43 $\pm$ 0.02 0.39 $\pm$ 0.02 - 6.79 $\pm$ 0.24 6.13 $\pm$ 0.09 7.70 $\pm$ 0.05 HOPSE-M 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00 0.32 $\pm$ 0.00 0.47 $\pm$ 0.03 0.35$\pm$ 0.01 5.13 $\pm$ 0.14 5.60 $\pm$ 0.55 5.54 $\pm$ 0.05 HOPSE-G 0.02 $\pm$ 0.00 0.10 $\pm$ 0.02 0.31 $\pm$ 0.00 0.37 $\pm$ 0.02 0.28 $\pm$ 0.01 4.48 $\pm$ 0.14 4.23 $\pm$ 0.12 4.48 $\pm$ 0.11 Cell CCCN 0.11 $\pm$ 0.00 0.32 $\pm$ 0.01 1.33 $\pm$ 0.02 1.33 $\pm$ 0.02 6.14 $\pm$ 0.04 - - - CWN 0.10 $\pm$ 0.00 0.37 $\pm$ 0.02 1.38 $\pm$ 0.03 1.37 $\pm$ 0.01 5.86 $\pm$ 0.05 - - - HOPSE-M 0.02 $\pm$ 0.00 0.13 $\pm$ 0.01 0.34 $\pm$ 0.00 0.33 $\pm$ 0.01 0.47 $\pm$ 0.02 - - - HOPSE-G 0.02 $\pm$ 0.00 0.09 $\pm$ 0.00 0.40 $\pm$ 0.02 0.37 $\pm$ 0.00 0.32 $\pm$ 0.01 - - -
F.5 Ablation on neighborhood set ${\mathcal{N}}_{C}$
To gain more insight into how the choice of neighborhoods ${\mathcal{N}}_{C}$ impacts the performance of the HOPSE framework, this section shows an ablation over the considered sets of neighborhoods. The neighborhood sets listed in the tables are defined in Table 5. Each configuration differs in the number and type of neighborhoods considered, directly influencing computational complexity and model expressivity. Specifically, the Adj family explores adjacency-based neighborhoods of increasing order: Adj-1 uses a single neighborhood, while Adj-2 and Adj-3 consider two and three neighborhoods, respectively. The Inc configurations mix the base adjacency (${\mathcal{A}}_{0,1}$) and incidence, with Inc-1 and Inc-2 using three neighborhoods, and Inc-3 extending to five. The Mix configurations, Mix-1 and Mix-2, are the biggest, combining a total of eight and six neighborhoods, respectively. Mix-1 uses all previous adjacencies and incidences combined while Mix-2 includes more complex neighborhoods. One of such is ${\mathcal{A}}_{0,2}$, which means the adjacencies of 00-cells with $2$-cells through $1$-cells. Effectively, this means that the structure of neighborhoods between non-adjacent ranks (i.e., a difference of $1$ in the rank number) is considered. As will be seen later in the section, considering these non-standard neighborhoods does account for a difference in performance for the more complex benchmarks.
@@ F.5 Ablation on neighborhood set ${\mathcal{N}}_{C}$
Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1}$
@@
F.5.1 Analysis of performance across neighborhoods
Analysis of performance for different domains.
The results in Table 6 and Table 7 demonstrate insightful patterns in how neighborhood selection impacts performance depending on domain type and dataset complexity. Comparing performance between HOPSE-M and HOPSE-G variants across different neighborhood configurations further suggests that the pre-training strategy employed in GPSE provides rich neighborhood information, particularly for the molecular datasets, as the GPSE encoder has been pre-trained on molecular data. This is most evident in the cell domain results for ZINC, where HOPSE-G with Mix-2 achieves a 16% improvement compared to the best HOPSE-M configuration.
@@ F.5.1 Analysis of performance across neighborhoods
Table 6: Performance per neighborhood on simplicial. Results are shown as mean $\pm$ standard deviation. The best result per neighborhood is in bold and results within one standard deviation of the best per neighborhood are highlighted in blue. The performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other. As datasets grow bigger and more complex, such as ZINC and the MANTRA benchmarks, the neighborhood choice starts to matter. The results show that for simplicial models, the most informative neighborhoods for ZINC are the extended groups of adjacencies Adj-2 and Adj-3. However, for the MANTRA tasks, the most important neighborhood is Mix-2

### v2

@@ Appendix F Experimental configuration
F.1 Datasets
The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^6] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [Morris et al., 2020]. For regression, the ZINC dataset is used [Irwin et al., 2012]. Following standard practice in the TDL literature [Telyatnikov et al., 2024], liftings are applied to transform these graph datasets into higher-order domains. Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
Additionally, the MANTRA Ballester et al. [2025] dataset comprises different triangulations of manifolds. A triangulation of a manifold $\mathcal{M}$} is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between $\mathcal{M}$} and the geometric realization of $K$. Concretely, triangulations encode higher-order topological information that the underlying graph, or 0-skeleton, does not encode. Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions. This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties. Thus, it aims to assess how much the model encodes topological information beyond the base graph. In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones. No other datasets are included, such as the ones proposed in [Eitan et al., 2024], as they are subsumed by MANTRA. Since MANTRA contains a wider variety of homeomorphic features, categorizing them by homeomorphism class (NAME) or their Betti numbers is more general and complex.
Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al. [2025] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number. The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT. Additionally the task of $\beta_{0}$ is excluded as all graphs are connected thus rendering the label as $1$ for all cases.
F.2 Training and evaluation.
Each dataset splits into three parts, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively. The exception is ZINC, for which the predefined splits are used [Irwin et al., 2012]. The results are the average over 5 runs using different seeds.
@@ F.1 Datasets
Discrepancy in ZINC Results and other datasets
The discrepancies in our baseline results from common literature arise from our deliberate choice to use the TopoBench Telyatnikov et al. [2024] suite, a framework designed for standardized and fair comparison. Our goal was to evaluate all methods on an even playing field to rigorously assess the performance-efficiency trade-off. This standardization introduces several key differences compared to prior publications [Corso et al., 2020, Dwivedi et al., 2023, Jin et al., 2024, Bouritsas et al., 2022]:
  • Experimental protocol: A single 50/25/25 train/validation/test split and a constrained hyperparameter search are used for all models (please refer to in Section˜F.3). This differs from the k-fold cross-validation and extensive tuning found in other works.
  • Standardized components: TopoBench uses common encoders and readouts to fairly evaluate the core “backbone” of each model. This contrasts with studies using domain-specific knowledge, such as directly encoding molecular rings.
  • Feature handling: To ensure a fair comparison, TopoBench sometimes disregards certain features (e.g., edge attributes in ZINC or MUTAG) in favour of standardization, which impacts the final reported performance.
F.3 Hyperparameter search
@@
  • Model hidden dimensions: {64, 128, 256 }
  • Encoder dropout: {0, 0.25, 0.5}
@@
Appendix G Complete model performance and runtimes
Table 3: Performance on benchmark datasets.
@@ Appendix G Complete model performance and runtimes
Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC-12K ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M (Our) 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G (Our) 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M (Our) 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.24 $\pm$ 0.01 - - - - HOPSE-G (Our) 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
Table 4: Total end-to-end training runtime in seconds.
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC-12K NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M (Ours) 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G (Ours) 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M (Ours) 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G (Ours) 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
Appendix H Performance ablation across different neighborhoods
The choice of neighborhood functions set ${\mathcal{N}}_{C}$ is a crucial and non-trivial choice for the performance of HOPSE. This section contains the performance of HOPSE for each choice of neighborhood encoded by a distinct name, as reported in Table˜5.
@@ Appendix H Performance ablation across different neighborhoods
Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
@@
For small and simple datasets (MUTAG, PROTEINS, NCI1, NCI109), the performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other. As datasets grow bigger and more complex, such as ZINC-12K and the MANTRA, the neighborhood choice starts to matter. The results show that for simplicial models, the best performing neighborhoods for ZINC-12K are the extended groups of adjacencies Adj-2 and Adj-3. However, for the MANTRA tasks, the most important neighborhood is Mix-2.Note that the per neighborhood ablation results differ from the main table due to some neighborhoods having a higher validation which does not lead to ahigher test.These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in the topological learning frameworks. While simpler configurations may suffice for some graph benchmarks, an effective strategy is to consider a greater neighbourhood set ${\mathcal{N}}_{C}$ to capture the full expressive power of topological structures.
Table 6: Performance ablation of $\mathcal{N}_{S}$ on simplicial domain.